the sum of two numbers is 23.twice the larger number is 4 more than 4 times the smaller.what are the numbers?
Answer by praseenakos@yahoo.com(507)

QUESTION:


the sum of two numbers is 23.twice the larger number is 4 more than 4 times the smaller.what are the numbers?


ANSWERS:

Assume that the larger number is 'x' and the smaller number is 'y'

Given that their sum is = 23



==> x + y = 23 --------------------(1)

Twice the largerr number cam be expressed as '2x'

four times the smaller number can be expressed as, '4x'


twice the larger number is 4 more than 4 times the smaller can be expressed as,


==> 2x = 4y + 4


==> 2x - 4y = 4 -----------------(2)



Multiply (1) throughout with 4



==> 4(x + y) = 4 * 23


==> 4x + 4y = 92 ---------------(3)

Now add (2) and (3)

2x - 4y = 4 -----------------(2)
4x + 4y = 92 ---------------(3)


==> 2x - 4y + 4x + 4y = 4 + 92


==> 6x = 96


Divide throughout by 6


==> 6x/6 = 96/6


==> x = 16


Now substitute this value of x in the given condition, x + y = 23


==> 16 + y = 23

Subtract 16 from both sides


==> 16 + y - 16 = 23 - 16


==> y = 7


So the larger number is 16and the smaller number is 7.



Hope you found the explanation useful.


Regards.


Praseena,